A geometric progression has first term and common ratio

i)                  Find the set of values of x for which geometric progression has a sum to infinity.

  • ii)                Find the value of x for which the sum to infinity of geometric progression 
1
Expert's answer
2021-07-06T14:32:22-0400

I haven't got any information on first term and common ratio of the geometric progression. I assume that the first term is "a\ne 0" and common ratio is x.

Then n-th term is "ax^{n-1}" and sum of the first n term is

"s_n=a_1+\dots+a_n=a+ax+\dots+ax^{n-1}=a\frac{x^n-1}{x-1}=\frac{a}{1-x}-\frac{a}{1-x}x^n" ("x\ne 1")

"s_n" tends to a finite limit as "n\to+\infty" if and only if "x^n" tends to a finite limit as "n\to+\infty", i.e., if and only if "|x|<1". With this condition we have "\lim\limits_{n\to+\infty}x^n=0" and "\sum\limits_{n=1}^{+\infty}ax^{n-1}=\lim\limits_{n\to+\infty}s_n=\frac{a}{1-x}".

In the case x=1 we have "a_n=ax^{n-1}=a" and "s_n=an" diverges to infinity.

Answer. (i) "|x|<1"; (ii) "\sum\limits_{n=1}^{+\infty}ax^{n-1}=\frac{a}{1-x}".